The balance sheet for Levy Corp. is shown here in market value terms. There are 6,000 shares of stock outstanding. Market Value Balance Sheet Cash $ 44,900 Equity $ 474,900 Fixed assets 430,000 Total $ 474,900 Total $ 474,900 Instead of a dividend of $1.80 per share, the company has announced a share repurchase of $10,800 worth of stock. How many shares will be outstanding after the repurchase? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Shares outstanding What will the price per share be after the repurchase? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

New stock price $?

Shares outstanding?
Answer #1

share price=474900/6000=79.15

How many shares will be outstanding after the repurchase

=6000-(10800/79.15)

=5863.55

What will the price per share be after the repurchase

=(474900-10800)/5863.55

=79.15

the above is answer..